Let 𝑃 be the midpoint of the line segment joining the points 𝐴(π‘Ž + 𝑏, 𝑏) and 𝐡(π‘Ž βˆ’ 𝑏, π‘Ž + 𝑏). Find the slope of the line passing through 𝑃 and 𝑄 (𝑏, βˆ’π‘Ž/2) .


1
Expert's answer
2022-02-07T10:17:24-0500

It follows that for the point "P(x,y)" we have that "x=\frac{a+b+a-b}2=a,\ y=\frac{b+a+b}2=\frac{a}2+b." Therefore, the slope of the line passing through "P(a,\frac{a}2+b)" and "𝑄 (𝑏, −\frac{𝑎}2)" is equal to

"\frac{\frac{a}2+b-(-\frac{a}2)}{a-b}=\frac{a+b}{a-b}."